OpenStudy (perl):

In a class of 30 students , 20 of them are boys . If two students are chosen at random, what is the probability they are different genders and are born on the same day of the year

OpenStudy (goformit100):

Take Probabalilty conception in use.

OpenStudy (perl):

P( two students have different gender & born on same day of year) = P( (BG or GB) & born same day of year)

OpenStudy (perl):

BG = choosing a boy , then choosing a girl

OpenStudy (perl):

P( (BG or GB) & born same day of year) = P( BG or GB) * P ( two students born same day of year) ?

OpenStudy (perl):

= ( 2*20/30 * 10/29) ( 1/365) ?

OpenStudy (perl):

it seems to me that they are independent

terenzreignz (terenzreignz):

gender does seem independent of birthdate

OpenStudy (perl):

did i do this wrong?

terenzreignz (terenzreignz):

I don't think so. But I'm not the best person to consult when statistics and probability is concerned... you'd best get a second opinion (or a third, if I'm the second)

OpenStudy (perl):

with general multiplication rule we have P( BG or GB) * P ( two students born same day of year | BG or GB)

OpenStudy (kropot72):

Whether a boy is chosen first or a girl is chosen first the probability of different genders being chosen is \[P(different\ genders)=\frac{20\times 10}{30\times 29}\] The required probability is \[\frac{20\times 10}{30\times 29\times365}\]

OpenStudy (perl):

but I got BG or GB

OpenStudy (perl):

so should be times 2

OpenStudy (perl):

you can pick a boy first, then girl, or you can pick a girl first then boy

OpenStudy (kropot72):

My bad. You are correct. There are two ways of choosing different genders, each way with the same probability. The probability of choosing a boy first is 20/30. The probability of choosing a girl next is 10/29. The probability of choosing a girl first is 10/30. The probability of choosing a boy next is 20/29. \[\frac{20\times 10}{30\times 29}=\frac{10\times 20}{30\times 29}\]